The dissociation constant of a weak acid HA and weak base BOH are 2 x 10-5 and 5 x 10-6 respectively. The equilibrium constant for the neutralization reaction of the two is (ignore hydrolysis of resulting salt)
 
(Ans: 1.0 x 10-10 )

Grade:12th pass

1 Answers

Pooja kuril
6 years ago
Weak acid HA ,Ka=2×10-5 and weak base BOH,KB=5×10-6Kh=Ka×KbKh=Keq=Ka×KbKeq=(2×10-5)(5×10-6)Keq=10×10-11Keq=1×10-10